THEORETICAL INVESTIGATION OF THE M^{+}-RG_{2} (M = ALKALINE EARTH METAL; RG = RARE GAS) COMPLEXES

\underline{ADRIAN M. GARDNER}, RICHARD J. PLOWRIGHT, JACK GRANEEK, TIMO-THY G. WRIGHT, School of Chemistry, University of Nottingham, University Park, Nottingham, NG7 2RD, UK; and W. H. BRECKENRIDGE, Department of Chemistry, University of Utah, 315 South 1400 East, Salt Lake City, Utah 84112, USA.

Metal cation rare gas complexes provide an expectedly simple system with which to investigate intermolecular interactions. Despite this, we have previously found the M^{+}-RG (M = alkaline earth metal) complexes to very complicated systems, with the complexes of the heavier rare gases displaying surprisingly large degrees of chemical character.^{a}^{,}^{b}^{,}^{c} Here we extend these studies by examining the nature of these interactions with increasing degrees of solvation through investigating the M^{+}-RG_{2} complexes using high level ab initio techniques. Intriguing trends in the geometries and dissociation energies of these complexes have been observed and are rationalized.

aA. M. Gardner, C. D. Withers, J. B. Graneek, T. G. Wright, L. A. Viehland and W. H. Breckenridge, J. Phys. Chem. A, 2000, 114, 7631.

bA. M. Gardner, C. D. Withers, T. G. Wright, K. I. Kaplan, C. Y. N. Chapman, L. A. Viehland, E. P. F. Lee and W. H. Breckenridge, J. Chem. Phys., 2010, 132, 054302.

cM. F. McGuirk, L. A. Viehland, E. P. F. Lee, W. H. Breckenridge, C. D. Withers, A. M. Gardner, R. J. Plowright and T. G. Wright, J. Chem. Phys., 2009, 130, 194305.